OpenStudy (anonymous):

How to do you solve exponential functions? I understand how to make the table of values and graph, but do not understand how to simplify the expressions involving add, multiply, subtract the exponents. Can someone please help me?

OpenStudy (anonymous):

if you can give an example that would help

OpenStudy (anonymous):

http://www.purplemath.com/modules/simpexpo.htm

OpenStudy (anonymous):

\[7\sqrt[5]{5}/7\sqrt[2]{5}\]

OpenStudy (phi):

The easiest way to handle these is realize there are two ways to write the same thing \[\sqrt[a]{x}= x^{\frac{1}{a}}\]

OpenStudy (anonymous):

\[(y \sqrt[2]{2})^{2}\]

OpenStudy (phi):

so we could write your problem as \[ \frac{7\sqrt[5]{5}}{7\sqrt[2]{5} }= \frac{5^{\frac{1}{5}}}{5^{\frac{1}{2}}}\] \[ 5^{(\frac{1}{5}-\frac{1}{2})} = 5^{-\frac{3}{10}}\]

OpenStudy (phi):

for \[ ( y \sqrt[2]{2})^{2} = (y\cdot 2^{\frac{1}{2}})^2= y^2\cdot 2^{\frac{1}{2}\cdot 2}=2y^2\]

OpenStudy (phi):

btw, we can change 5^(-3/10) back into radical form. First, the minus sign means flip it, put the number in the denominator. \[ 5^{-\frac{3}{10}} = \frac{1}{5^{\frac{3}{10}} } \] the 1/10 can be written as a radical \[\frac{1}{5^{\frac{3}{10}} }= \frac{1}{\sqrt[10]{5^3}}\]

OpenStudy (phi):

which problem?

OpenStudy (anonymous):

\[(y \sqrt[2]{2})^{3}\]

OpenStudy (phi):

First, can you rewrite the sqrt(2) as 2^(1/2)?

OpenStudy (anonymous):

All of the above if possible. Sorry

OpenStudy (phi):

can you rewrite the sqrt(2) as 2^(1/2)?

OpenStudy (phi):

Does that make sense? We are using an idea someone came up with. They proved that radicals could be rewritten as a fractional power.

OpenStudy (phi):

If you don't understand you might watch Khan's video http://www.khanacademy.org/math/algebra/exponents-radicals/v/fractional-exponent-expressions-2 when you have time.

OpenStudy (phi):

But if you believe it, we can continue \[( y \sqrt[2]{2})^{3} = (y\cdot 2^{\frac{1}{2}})^3\]